OpenStudy (anonymous):

Write an equation in standard form for a line with a slope of -1 passing through C(2,1)

OpenStudy (texaschic101):

Standard form : Ax + By = C y = mx + b slope(m) = -1 (2,1)...x = 2 and y = 1 now we sub 1 = -1(2) + b 1 = -2 + b 1 + 2 = b 3 = b equation is : y = -x + 3, but we need it in standard form y = -x + 3 -- add x to both sides x + y = 3 <== standard form

OpenStudy (anonymous):

how did you get the equation?

OpenStudy (anonymous):

wait nvm. thank you so much for ur help